# Provable Robustness in Multimodal Large Language Models via Feature Space Smoothing

### v1

@@ 4.1 Noise purifier
which encourages feature consistency between the purified and clean representations. Practically, we adopt the ImageNet-pretrained guided-diffusion [7] as our purifier $\mathcal{P}$ and perform one-step diffusion for denoising in our main experiments. We then fine-tune this model on the dataset $\mathcal{D}$ using the purifier loss $\mathcal{L}_{\mathcal{P}}$, which is:
$\vskip-2.84526pt\mathcal{L}_{\mathcal{P}}=l_{\mathrm{diff}}+\lambda_{1}l^{\mathcal{P}}_{\mathrm{rb}}+\lambda_{2}l_{mse},$ (9)
where $l_{\mathrm{diff}}$ denotes the original diffusion loss, and $\lambda_{1}$ and $\lambda_{2}$ are the weighting coefficients. More details of the fine-tuning process on $\mathcal{P}$ can be found in the supplementary material, Section S.2.
@@ 4.2 Residual smoothness mapper
$\vskip-2.84526pt\mathcal{L}_{\mathcal{M}}=l^{\mathcal{M}}_{\mathrm{rb}}+\lambda_{3}l_{\mathrm{stats}}+\lambda_{4}l_{\mathrm{id}},$ (13)
where $\lambda_{3}$ and $\lambda_{4}$ are the weighting coefficients. More details of the training process on the residual smoothness mapper can be found in the supplementary material, Section S.2.
4.3 Further discussion on PSM
Certified robustness for the encoder $f_{e}$ with PSM. Let $f^{\prime}_{e}$ denote the feature encoder integrated with the proposed PSM. Then the forward process can be formulated as:
@@ 4.3 Further discussion on PSM
Where $\hat{S}^{\prime}(\bm{x})\in[0,1]$. We can prove that the Lipschitz property derived in Lemma 1 still holds, and the theoretical bound in Section 3 remains valid. Then, utilizing Theorem 1 and Corollary 1, we can derive the certified lower bound on $\mathrm{Cos}(\hat{f}^{\prime}_{e}(\bm{x^{\prime}}),f_{e}(\bm{x}))$ for any adversarial input $\bm{x}$$^{\prime}$.
Algorithm 1 Training algorithm of PSM
1:Input: dataset $\mathcal{D}$, feature encoder $f_{e}$, purifier $\mathcal{P}$, mapper $\mathcal{M}$, sampling number $n_{0}$, noise standard deviation $\sigma$, and loss weights $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$.
2:Output: trained purifier $\mathcal{P}$ and mapper $\mathcal{M}$.
3:// Stage 1: Fine-tune the purifier $\mathcal{P}$
4:for each batch ${\bm{x}}\sim\mathcal{D}$ do
5:  ${\bm{z}}=f_{e}({\bm{x}})$
6:  for $i=1,\dots,n_{0}$ do
7:    Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
8:    $\tilde{{\bm{x}}}^{i}=\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$, and $\tilde{{\bm{z}}}^{i}=f_{e}(\tilde{{\bm{x}}}^{i})$
9:  Compute $l^{\mathcal{P}}_{\mathrm{rb}}$ and $l_{\mathrm{mse}}$ using Equations 78.
11:// Stage 2: Train the smoothness mapper $\mathcal{M}$
12:for each batch ${\bm{x}}\sim\mathcal{D}$ do
13:  ${\bm{z}}=f_{e}({\bm{x}})$
14:  for $i=1,\dots,n_{0}$ do
15:    Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
16:    $\tilde{{\bm{x}}}^{i}=\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$, and $\tilde{{\bm{z}}}^{i}_{m}=f_{e}(\tilde{{\bm{x}}}^{i})+\mathcal{M}(\tilde{{\bm{z}}}^{i},\sigma)$
17:  Compute $l^{\mathcal{M}}_{\mathrm{rb}}$, $l_{\mathrm{id}}$ and $l_{\mathrm{stats}}$ using Equations 1112.
18:  Update $\mathcal{M}$ by gradient descent on $\nabla_{\mathcal{M}}\mathcal{L}_{\mathcal{M}}$
Table 1: Experimental results on adversarial robustness of different defense methods and MLLMs on image captioning tasks. Values in parentheses denote the average textual similarity measured by GPTScore. The overall best results are shown in bold, and the best results without smoothing are underlined, highlighting the significant performance gain introduced by our proposed FS-PSM.
Model Method M-Attack [23] FOA [14] AttackVLM [51]
[…]
Smoothed org. 0.650 82% (0.694) 2% (0.028) 0.652 87% (0.735) 1% (0.021) 0.660 89%(0.727) 2% (0.018)
FARE [37] 0.588 44% (0.409) 24% (0.097) 0.504 19% (0.197) 51% (0.222) 0.499 18%(0.190) 44%(0.221)
Smoothed FARE 0.716 59% (0.501) 14% (0.069) 0.656 35% (0.327) 29% (0.153) 0.729 60% (0.503) 11% (0.061)
TeCoA [29] 0.720 51% (0.458) 16% (0.081) 0.674 21% (0.236) 37% (0.187) 0.606 17% (0.179) 43% (0.218)
Smoothed TeCoA 0.805 59% (0.498) 14% (0.056) 0.789 39% (0.364) 20% (0.094) 0.755 42%(0.358) 18% (0.104)
Open Flamingo -9B Org. 0.351 1%(0.089) 86%(0.583) 0.347 1%(0.091) 87%(0.569) 0.442 16%0.222) 59%(0.379)
Smoothed org. 0.673 73%(0.655) 1%(0.063) 0.672 69%(0.623) 0%(0.061) 0.703 73%(0.687) 0%(0.055)
FARE [37] 0.588 36%(0.407) 17%(0.145) 0.504 20%(0.249) 40%(0.240) 0.499 15%(0.243) 40%(0.247)
Smoothed FARE 0.716 55%(0.529) 9%(0.088) 0.656 40%(0.414) 17%(0.125) 0.729 50%(0.488) 4%(0.089)
TeCoA [29] 0.720 51%(0.472) 13%(0.116) 0.674 21%(0.240) 30%(0.210) 0.606 19%(0.244) 34%(0.225)
Smoothed TeCoA 0.805 56%(0.518) 5%(0.090) 0.789 50%(0.454) 8%(0.095) 0.755 38%(0.398) 12%(0.126)
Training algorithm and dataset construction. The training procedure of PSM is summarized in Algorithm 1, where the purifier $\mathcal{P}$ and the smoothness mapper $\mathcal{M}$ are trained sequentially via a two-stage manner. The expectation over Gaussian perturbations is approximated by Monte Carlo sampling with $n_{0}$ samples drawn from $\mathcal{N}(0,\sigma^{2}I)$, where we set $n_{0}=8$ in our training to balance efficiency and estimation accuracy. Notably, all modules can be trained in a self-supervised fashion, without relying on task-specific labels. Thus, to promote generalization across diverse visual domains and ensure compatibility with various MLLMs’ downstream tasks, we collect the training dataset $\mathcal{D}$ comprising 5,000 images from heterogeneous domains, including medical, cartoon, and natural images.
@@
CLIP-B16 0.25 0.828 0.623 0.313 -0.06 / 0.31
@@
CLIP-B16 0.50 0.717 0.586 0.439 0.253 -0.132 0.33
[…]
Certified robustness bound of FS-PSM. To evaluate the effectiveness of the proposed PSM module, we adopt CLIP-B16 [33] as the vanilla feature encoder $f_{e}$ and assess the certified robustness bound of both the original smoothed encoder $\hat{f}_{e}$ and its PSM-enhanced counterpart $\hat{f}^{\prime}_{e}$ under the FS. We report both the certified FCSB under different adversarial constraints $\epsilon$ and the certified radius $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$. The results are shown in Table 2. For implementation, we randomly sample 100 images from the ImageNet dataset [6] and approximate the Gaussian expectation using n=1,000n=1{,}000 Monte Carlo samples. As shown in Table 2, integrating PSM consistently enhances both the certified FCSB across a range of perturbation magnitudes and the average certified radius $\mathcal{R}$, indicating substantial improvements in certified robustness. To further assess the practical protection of our FS-PSM, we then conduct extensive experiments across multiple MLLMs and downstream tasks in the next section, evaluating their performance before and after smoothing under various strong white-box attacks.
5 Experiments
@@ 5.1 Experimental setup
Evaluated models and tasks. Since the proposed FS requires access to the forward feature computation process of MLLMs, we primarily validate its effectiveness on open-sourced MLLMs, including LLaVA-V1.5-7B [24, 25] and OpenFlamingo9B [1]. We comprehensively assess the performance of plugging the FS-PSM under adversarial conditions across multiple downstream tasks, including:
  • Image captioning: Following the setup in [23, 14], we randomly take 100 images from the NIPS 2017 Adversarial Attacks and Defenses Competition dataset 5.1 and ask the model to caption the image.
  • Image Classification: We randomly sample 500 images from 10 classes in the ImageNet dataset [6] and ask the model to classify the input.
  • Visual Question Answering (VQA): We utilize 100 image-question pairs from the ScienceQA dataset [26].
@@
Meanwhile, we also evaluate our FS-PSM on the original CLIP-L14 [33] on image classification for the generality.
@@
Model Method M-Attack [23] FOA [14] AttackVLM [51]
@@
Smoothed org. 0.681 90.4% 0.6% 0.697 88.8% 0.4% 0.700 88.4% 0.2%
FARE [37] 0.574 55.4% 14.6% 0.521 24.8% 39.8% 0.508 25.8% 45.0%
Smoothed FARE 0.859 75.6% 5.2% 0.712 72.0% 5.6% 0.699 74.0% 6.2%
TeCoA [29] 0.731 67.4% 2% 0.626 29.0% 20.0% 0.592 33.6% 17.8%
Smoothed TeCoA 0.819 74.4% 0.4% 0.727 57.2% 3.0% 0.711 57.4% 2.0%
CLIP -L14 Org. 0.427 18.0% 37.2% 0.437 10.0 % 52.4 % 0.458 16.0% 48.4%
Smoothed org. 0.681 92.4% 0.4% 0.697 92.0% 0.4% 0.700 92.0% 0.4%
FARE [37] 0.574 70.4% 5.2% 0.521 42.0% 14.4% 0.508 35.6% 24.4%
Smoothed FARE 0.859 85.2% 0.8% 0.712 81.6% 0% 0.699 82.4% 0%
TeCoA [29] 0.731 81.2% 2.4% 0.626 39.2% 20.2% 0.592 35.8% 24.4%
Smoothed TeCoA 0.819 88.4% 1.6% 0.727 76.0% 4.8% 0.711 73.2% 6.0%
The threat model. To comprehensively assess the robustness under strong adversaries, we employ three SOTA adversarial attacks specifically designed for MLLMs, named AttackVLM [51], M-Attack [23], and FOA [14] using the white-box setting. All attacks are implemented following their original best configurations, with the adversarial perturbation budget $\epsilon$ set to $\|\epsilon\|_{\infty}=16/255$. We also consider a stronger bound with $\|\epsilon\|_{\infty}=32/255$ for a stress test.
Compared defense methods. As this work pioneers the research for MLLMs’ certified robustness, we primarily compare our defense with adversarial training based approaches, namely FARE [37] and TeCoA [29]. Both FARE and TeCoA adopt adversarial training to obtain robust feature encoders that can be directly integrated into models such as LLaVA-1.5-7B and OpenFlamingo-9B. To ensure a fair and consistent comparison, we obtain all adversarially trained feature encoders from their official repositories, without any modification.
Implementation details. For practical inference efficiency, we set the number of samples to $n_{0}=4$ for smoothing, which provides a favorable trade-off between robustness and runtime. We train three independent PSM modules, utilizing the vanilla feature encoders from models including LLaVA-1.5-7B, OpenFlamingo-9B, and CLIP-L14. To further assess cross-model generalization, we evaluate the PSM trained on a vanilla encoder by directly applying it to adversarially trained encoders FARE and TeCoA. For all feature encoders, we set hyperparameters $\lambda_{1},~\lambda_{2},~\lambda_{4}=0.25$, $\lambda_{3}=100$, and $\sigma=0.25$. In our tables, the term ”smoothed” denotes the process of smoothing the encoder via FS and enhancing it with PSM.
Evaluation metrics. We mainly report: the Feature Cosine Similarity (FCS), the Accuracy (ACC), and the Attack Success Rate (ASR). Specifically, FCS measures the cosine similarity between clean and adversarial features extracted by the same feature encoder, reflecting its feature-space robustness. ACC denotes the proportion of correctly completed tasks, while ASR indicates the proportion of cases where the model is successfully manipulated to produce the adversarially targeted outputs. For image classification and VQA tasks, ACC and ASR are determined by whether the MLLM outputs match the correct or adversarial targets.
For image captioning, following [23], we adopt the LLM-as-a-judge protocol [52] to evaluate ACC and ASR. Specifically, we first generate the clean and adversarially targeted captions by feeding the clean and targeted inputs into the vanilla MLLM (e.g., the original LLaVA). We then obtain the predicted caption by feeding the adversarially perturbed input into the tested model (e.g., Smoothed LLaVA). The textual similarity is computed using GPTScore [9], where a task is considered successful if the GPTScore between the predicted and clean captions is $\mathbf{\geq 0.5}$, and an attack is considered successful if the GPTScore between the predicted and adversarially targeted captions is $\mathbf{\geq 0.3}$.
@@
Model Method M-Attack [23] FOA [14] AttackVLM [51]
@@
Smoothed org. 0.732 47% 0% 0.676 43% 0% 0.765 43% 1%
FARE [37] 0.657 47% 5% 0.590 32% 0% 0.550 38% 7%
Smoothed FARE 0.859 48% 1% 0.768 38% 0% 0.818 39% 0%
TeCoA [29] 0.788 31% 2% 0.748 31% 2% 0.667 31% 1%
Smoothed TeCoA 0.917 32% 1% 0.858 34% 1% 0.862 29% 1%

### v2

@@ 4.1 Gaussian Denoiser
which encourages feature consistency between the purified and clean representations. We then fine-tune this model on the dataset $\mathcal{D}$ using the purifier loss $\mathcal{L}_{\mathcal{P}}$, which is:
$\vskip-2.84526pt\mathcal{L}_{\mathcal{P}}=\ l^{\mathcal{P}}_{\mathrm{rb}}+\lambda_{1}l_{mse},$ (9)
where $\lambda_{1}$ is the weighting coefficient. More details of the fine-tuning process on $\mathcal{P}$ can be found in the Appendix A.2.1.
@@ 4.2 Residual Smoothness Mapper
$\vskip-2.84526pt\mathcal{L}_{\mathcal{M}}=l^{\mathcal{M}}_{\mathrm{rb}}+\lambda_{2}l_{\mathrm{stats}}+\lambda_{3}l_{\mathrm{id}},$ (13)
where $\lambda_{2}$ and $\lambda_{3}$ are the weighting coefficients. More details of the training process on the residual smoothness mapper can be found in the Appendix A.2.2.
Algorithm 1 Training algorithm of GSB
0:  Dataset $\mathcal{D}$, feature encoder $f_{e}$, Denoiser $\mathcal{P}$, mapper $\mathcal{M}$, sampling number $n_{0}$, noise std $\sigma$, loss weights $\lambda_{1},\lambda_{2},\lambda_{3}$.
1:  Stage 1: Train denoiser $\mathcal{P}$.
2:  for each batch ${\bm{x}}\sim\mathcal{D}$ do
3:  ${\bm{z}}\leftarrow f_{e}({\bm{x}})$
4:  for $i=1$ to $n_{0}$ do
5:   Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
6:   $\tilde{{\bm{x}}}^{i}\leftarrow\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$
7:   $\tilde{{\bm{z}}}^{i}\leftarrow f_{e}(\tilde{{\bm{x}}}^{i})$
8:  end for
9:  Compute $l^{\mathcal{P}}_{\mathrm{rb}}$ and $l_{\mathrm{mse}}$ using Eqs. 7 and 8.
11:  end for
12:  Stage 2: Train smoothness mapper $\mathcal{M}$.
13:  for each batch ${\bm{x}}\sim\mathcal{D}$ do
14:  ${\bm{z}}\leftarrow f_{e}({\bm{x}})$
15:  for $i=1$ to $n_{0}$ do
16:   Sample $\bm{\varepsilon}^{i}\sim\mathcal{N}(0,\sigma^{2}I)$
17:   $\tilde{{\bm{x}}}^{i}\leftarrow\mathcal{P}({\bm{x}}+\bm{\varepsilon}^{i})$
18:   $\tilde{{\bm{z}}}^{i}\leftarrow f_{e}(\tilde{{\bm{x}}}^{i})$
19:   $\tilde{{\bm{z}}}^{i}_{m}\leftarrow\tilde{{\bm{z}}}^{i}+\mathcal{M}(\tilde{{\bm{z}}}^{i},\sigma)$
20:  end for
21:  Compute $l^{\mathcal{M}}_{\mathrm{rb}}$, $l_{\mathrm{id}}$, and $l_{\mathrm{stats}}$ using Eqs. 11 and 12.
22:  Update $\mathcal{M}$ by gradient descent on $\nabla_{\mathcal{M}}\mathcal{L}_{\mathcal{M}}$
23:  end for
4.3 Further Discussion on Plugging GSB
Certified robustness for the encoder $f_{e}$ with GSB. Let $f^{\prime}_{e}$ denote the feature encoder integrated with the proposed GSB. Then the forward process can be formulated as:
@@ 4.3 Further Discussion on Plugging GSB
Where $\hat{S}^{\prime}(\bm{x})\in[0,1]$. We can prove that the Lipschitz property derived in Lemma 3.1 still holds, and the theoretical bound in Section 3 remains valid. Then, utilizing Theorem 3.2 and Corollary 3.3, we can derive the certified lower bound on $\mathrm{Cos}(\hat{f}^{\prime}_{e}(\bm{x^{\prime}}),f_{e}(\bm{x}))$ for any adversarial input $\bm{x}$$^{\prime}$.
Training algorithm and dataset construction. The training procedure of GSB is summarized in Algorithm 1, where the denoiser $\mathcal{P}$ and the smoothness mapper $\mathcal{M}$ are trained sequentially via a two-stage manner. The expectation over Gaussian perturbations is approximated by Monte Carlo sampling with $n_{0}$ samples drawn from $\mathcal{N}(0,\sigma^{2}I)$, where we set $n_{0}=8$ in our training to balance efficiency and estimation accuracy.
5 Certified Robustness via FS
5.1 Feature-wise Certification
@@
CLIP-L14 0.25 0.828 0.623 0.313 -0.06 / 0.31
CLIP-L14+GSB 0.907 0.763 0.508 0.152 / 0.38
CLIP-L14 0.50 0.717 0.586 0.439 0.253 -0.132 0.33
CLIP-L14+GSB 0.924 0.873 0.799 0.698 0.409 0.69
To evaluate the effectiveness of the proposed FS framework with GSB, we adopt CLIP-L14 (Radford et al., 2021) as the vanilla feature encoder $f_{e}$ and assess the feature-wise certified robustness under the FS. We report both the certified FCSB under different adversarial constraints $\epsilon$ and the average certified radius $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$. The results are shown in Table 1. For implementation, we randomly sample 500 images from the ImageNet dataset (Deng et al., 2009) and approximate the Gaussian expectation using n=10,000n=10{,}000 Monte Carlo samples. As shown in Table 1, integrating GSB consistently enhances both the certified FCSB across a range of perturbation magnitudes and the average certified radius $\mathcal{R}$, indicating substantial improvements in certified robustness.
5.2 Prediction-wise Certification
Beyond the feature-wise certification, the guarantees provided by FS can be propagated to the prediction space by smoothing the deepest-layer features with a solvable classification head, which enables a closed-form translation from feature-space certification to prediction-level robustness.
To this end, we adopt CLIP-L14 with GSB as the last-layer feature extractor for downstream classification. Within the FS framework, we employ a learnable prototype-based cosine classification head, where each class is represented by a unit-norm prototype, and predictions are determined by cosine similarity in the normalized feature space. This design admits a direct characterization of class decision boundaries, allowing the certified feature-wise cosine similarity guaranteed by FS to be converted into certified prediction bounds. We also compared the performance with Randomized Smoothing (RS) (Cohen et al., 2019), where we utilize the same CLIP-L14+GSB as the feature encoder and train a learnable linear classification head for a fair comparison.
In the experiments, we randomly sample 500 training instances from 10 ImageNet classes (Deng et al., 2009) to train both classification heads and evaluate them on an additional 500 test samples. For both FS and RS, the Gaussian expectation is approximated using n=10,000n=10{,}000 Monte Carlo samples. The comparative results in Table 2 demonstrate the effectiveness of the FS in providing certified prediction-level robustness. Further illustration and proof of prediction-wise certification via FS are provided in Appendix A.3.
Table 2: The comparison results of the RS and FS framework on prediction-wise certification.
Smoothing $\sigma$ Certified accuracy at different $\ell_{2}$ adv-bound $\varepsilon$ (%)
0.00 0.125 0.25 0.375 0.5 0.75 1 1.50
RS $0.25$ 93.4 88 86.2 81.0 75.4 66.5 / /
FS 94.2 92.4 90.0 88.0 85.8 72.0 / /
RS $0.50$ 91.2 86 84.4 81.4 77.0 71.6 67.0 49.2
FS 91 85.2 84.0 83.6 83.0 77.6 69.2 54.0
6 Experimental Results on MLLM Protection
While end-to-end robustness certification for MLLMs is appealing and practically meaningful, smoothing the entire model and certifying its autoregressive predictions with a solvable head are both computationally expensive and challenging. Alternatively, we thereby seek the feature-wise certification of the visual encoder (e.g., CLIP) in MLLMs, and empirically evaluate the resulting task-level robustness under strong white-box adversarial attacks, providing a practical compromise between certification feasibility and effectiveness. Notably, we find that this feature-wise certification offers two advantages: (i) efficiency, as smoothing the lightweight encoder $f_{e}$ incurs substantially lower computational overhead than smoothing the full model $\mathcal{F}$; and (ii) generality, where feature-representation level certification makes it applicable to various downstream tasks.
Table 3: Experimental results on adversarial robustness of different defense methods and MLLMs on image captioning tasks. Values in parentheses denote the average textual similarity measured by GPTScore. The overall best results are shown in bold, and the best results without smoothing are underlined, highlighting the significant performance gain introduced by our proposed FS.
Model Method M-Attack (Li et al., 2025b) FOA (Jia et al., 2025) AttackVLM (Zhao et al., 2023)
@@ 6 Experimental Results on MLLM Protection
Smoothed org. 0.588 63% (0.515) 3% (0.027) 0.587 65% (0.519) 6% (0.028) 0.598 71%(0.556) 4% (0.03)
FARE (Schlarmann et al., 2024) 0.588 44% (0.409) 24% (0.097) 0.504 19% (0.197) 51% (0.222) 0.499 18%(0.190) 44%(0.221)
Smoothed FARE 0.687 64% (0.541) 10% (0.053) 0.634 43% (0.391) 14% (0.065) 0.653 52% (0.443) 7% (0.045)
TeCoA (Mao et al., 2023) 0.720 51% (0.458) 16% (0.081) 0.674 21% (0.236) 37% (0.187) 0.606 17% (0.179) 43% (0.218)
Smoothed TeCoA 0.801 61% (0.510) 17% (0.081) 0.769 39% (0.364) 10% (0.0530) 0.720 38% (0.355) 14% (0.065)
Open Flamingo -9B Org. 0.351 1% (0.089) 86% (0.583) 0.347 1% (0.091) 87%(0.569) 0.442 16% (0.222) 59% (0.379)
Smoothed org. 0.592 59% (0.483) 3% (0.081) 0.588 57% (0.472) 1% (0.066) 0.703 61% (0.505) 0% (0.051)
FARE (Schlarmann et al., 2024) 0.588 36% (0.407) 17% (0.145) 0.504 20% (0.249) 40% (0.240) 0.499 15% (0.243) 40% (0.247)
Smoothed FARE 0.782 48% (0.462) 5% (0.062) 0.768 35% (0.354) 9% (0.102) 0.754 40% (0.404) 5% (0.086)
TeCoA (Mao et al., 2023) 0.720 51% (0.472) 13% (0.116) 0.674 21% (0.240) 30% (0.210) 0.606 19% (0.244) 34% (0.225)
Smoothed TeCoA 0.835 48% (0.464) 3% (0.067) 0.818 39% (0.371) 5% (0.088) 0.805 35%(0.341) 8%(0.079)
@@ 6.1 Experimental Setup
Evaluated models and tasks. Since the proposed FS requires access to the forward feature computation process of MLLMs, we primarily validate its effectiveness on open-sourced MLLMs, including LLaVA-V1.5-7B (Liu et al., 2023, 2024) and OpenFlamingo9B (Awadalla et al., 2023). We comprehensively assess the performance of plugging the FS under adversarial conditions across multiple downstream tasks, including:
  • Image captioning: Following (Li et al., 2025b; Jia et al., 2025), we randomly take 100 images from the NIPS 2017 Adversarial Attacks and Defenses Competition dataset 6.1 and ask the model to caption the image.
  • Image Classification: We randomly sample 500 images from 10 classes in the ImageNet dataset (Deng et al., 2009) and ask the model to classify the input.
  • Visual Question Answering (VQA): We utilize 100 image-question pairs from the ScienceQA dataset (Lu et al., 2022) and ask the model to select the answer.
@@
The threat model. To comprehensively assess the robustness under strong adversaries, we employ three SOTA adversarial attacks specifically designed for MLLMs, named AttackVLM (Zhao et al., 2023), M-Attack (Li et al., 2025b), and FOA (Jia et al., 2025) using the white-box setting, where attackers can fully access the feature extractor and GSB. All attacks are implemented following their original best configurations, with the adversarial perturbation budget $\epsilon$ set to $\|\epsilon\|_{\infty}=16/255$.
Compared defense methods. As robustness certification for MLLMs remains largely unexplored, we primarily compare our method against adversarial training–based defenses, specifically FARE (Schlarmann et al., 2024) and TeCoA (Mao et al., 2023). Both FARE and TeCoA adopt adversarial training to obtain robust feature encoders that can be directly integrated into models such as LLaVA-1.5-7B and OpenFlamingo-9B. To ensure a fair and consistent comparison, we obtain all adversarially trained feature encoders from their official repositories, without any modification.
Implementation details. For practical inference efficiency, we set the number of samples to $n_{0}=8$ for smoothing, ensuring a favorable trade-off between robustness and runtime. We train three independent GSB modules for different models, including the encoder of LLaVA-1.5-7B, the encoder of OpenFlamingo-9B, and CLIP-L14. To further assess cross-model generalization, we directly utilize the GSB trained on a vanilla encoder on adversarially trained encoders, including FARE and TeCoA. For all feature encoders, we set hyperparameters $\lambda_{1},~\lambda_{3}=0.25$, $\lambda_{2}=100$, and $\sigma=0.25$. In our tables, the term ”smoothed” denotes the process of smoothing the encoder via FS and enhancing it with GSB.
Evaluation metrics. We mainly report: the Feature Cosine Similarity (FCS), the Accuracy (ACC), and the Attack Success Rate (ASR). Specifically, FCS measures the cosine similarity between clean and adversarial features extracted by the same feature encoder, reflecting its feature-wise robustness. ACC denotes the proportion of correctly completed tasks, while ASR indicates the proportion of cases where the model is successfully manipulated to produce the adversarially targeted outputs. For image classification and VQA tasks, ACC and ASR are determined by whether the MLLM outputs match the correct or adversarial targets.
For image captioning, following (Li et al., 2025b), we adopt the LLM-as-a-judge protocol (Zheng et al., 2023) to evaluate ACC and ASR. Specifically, we first generate the clean and adversarially targeted captions by feeding the clean and targeted inputs into the vanilla MLLM (e.g., the original LLaVA). We then obtain the predicted caption by feeding the adversarially perturbed input into the tested model (e.g., Smoothed LLaVA). The textual similarity is computed using GPTScore (Fu et al., 2024), where a task is considered successful if the GPTScore between the predicted and clean captions is $\mathbf{\geq 0.5}$, and an attack is considered successful if the GPTScore between the predicted and adversarially targeted captions is $\mathbf{\geq 0.3}$.
@@
Model Method M-Attack (Li et al., 2025b) FOA (Jia et al., 2025) AttackVLM (Zhao et al., 2023)
@@
Smoothed org. 0.590 84.8% 0.2% 0.600 87.2% 0.2% 0.605 87.0% 0.2%
FARE (Schlarmann et al., 2024) 0.574 55.4% 14.6% 0.521 24.8% 39.8% 0.508 25.8% 45.0%
Smoothed FARE 0.695 75.6% 5.2% 0.681 66.2% 1% 0.671 68.0% 1%
TeCoA (Mao et al., 2023) 0.731 67.4% 2% 0.626 29.0% 20.0% 0.592 33.6% 17.8%
Smoothed TeCoA 0.790 70.6% 0.6% 0.714 58.8% 0.8% 0.712 56.4% 0.4%
@@
Model Method M-Attack (Li et al., 2025b) FOA (Jia et al., 2025) AttackVLM (Zhao et al., 2023)
@@
Smoothed org. 0.682 47% 0% 0.643 43% 0% 0.712 43% 1%
FARE (Schlarmann et al., 2024) 0.657 47% 5% 0.590 32% 0% 0.550 38% 7%
Smoothed FARE 0.861 48% 1% 0.788 38% 0% 0.825 39% 0%
TeCoA (Mao et al., 2023) 0.788 31% 2% 0.748 31% 2% 0.667 31% 1%
Smoothed TeCoA 0.925 32% 1% 0.880 34% 1% 0.879 29% 1%